OpenStudy (need.answers):

How would I graph this function: f(x) = 490 – 2/3x

OpenStudy (angelica):

I'm not sure but I think first you have to set values for x (i.e. -1, 0, 1, 2, 3) and substitute each value to the given f(x)

OpenStudy (angelica):

Then solve it to find the f(x) a.k.a the y. You might want to make a table so it would be easy. I hope this answers your question

hero (hero):

Step 1: Find the x and y intercepts. In terms of points they will look like this (x,0) and (0,y). (You have to find the values of x and y) Step 2: Plot the points on an xy plane. Step 3. Draw a straight line through the points.

OpenStudy (need.answers):

I don't know how to find these points you speak of.

OpenStudy (angelica):

The points are the values you use in your quadrants. It is in the positive and negative axes. Example X | Y -1| (substitute -1 in the f(x)) The result will be the points

OpenStudy (need.answers):

So, in f(x)=490-2/3x, I have to substitute -1, 0, 1, 2, etc, to find the points?

OpenStudy (angelica):

Yes

OpenStudy (need.answers):

So then.. f(-1)=490-2/3x would be a start to finding these points?

OpenStudy (angelica):

Yes, but you also shouldn't forget about the other x. The answer that you get from substituting will be the y-intercept. Which means that f(x) is equal to y. To plot the points you will be needing the value of x that you used for x-intercept and the y-intercept

OpenStudy (need.answers):

Right. I'm even more confused than when I began. Maybe I'll just get a job at McDonalds once I fail school. Have a medal anyhow for your effort, thanks!

OpenStudy (angelica):

Oh hey, that's not a good way of thinking. We can finish this. What are you confused about?

OpenStudy (need.answers):

Everything pretty much. I did get the art of substituting numbers for x out of the function, but after that I'm not sure what's going on.

OpenStudy (need.answers):

Part*

hero (hero):

What happens if you let x = 0, then solve for y as follows: y = 490 - (2/3)*(0)

hero (hero):

You only need two points two create a graph of a line btw. Through any two points exists exactly one line.

hero (hero):

f(x) is a linear function as you already know.

OpenStudy (angelica):

Oh okay. You substituted: f(x) = 490-2/3x with -1 for x. f(-1) = 490-2/3(-1) Then you solve that and you get the y Since there is 2 valuse needed to plot a point in the quadrants you'll need the x and y. To plot points, find -1 then the value of y

OpenStudy (angelica):

What @Hero said is right

OpenStudy (need.answers):

Okay, so substituting for x gives you the y?

OpenStudy (angelica):

Pretty much, yes

OpenStudy (need.answers):

So how do I then find the other point I need to graph the line.

OpenStudy (angelica):

You put another value for x

OpenStudy (need.answers):

I don't understand, if substituting for x gives you y, then why do it twice. I thought to create a line you need both x an y, not y and y.

OpenStudy (angelica):

You do it twice so you get two points. And as hero said you only need two to create a linear graph So if you have -1 and the y then you get 1 point. If you take 3 or whatever value you want for x and solve for y, then you get another.

OpenStudy (angelica):

They are partners, you see. There should be (x,y) for 1 point. The x and y are called intercepts

OpenStudy (need.answers):

Ohhh, I see. I may not fully comprehend but I think I can do it. Only a little more school left anyways, then I'll never use any of this again in my lifetime. Thanks once again!

OpenStudy (angelica):

No problem! I think you'd do great!